14. Consider a container with a volume of 22.4 L filled with a gas at 1.00 atm and 0˚C. What will be the new pressure if the temperature increases to 298K? (1.09)


1
Expert's answer
2021-04-23T07:30:31-0400

"\frac{P_1}{T_1} = \frac{P_2}{T_2} ( at-constant -volume)"

P1 = 1atm , T1 = 0⁰C = 273K , P2 = ? , T2 = 298K


"\frac{1atm}{273K} = \frac{P_2}{298K}"


"P_2 = \frac{1atm×298K}{273K} = 1.09 atm"